# Data Mixing Agent: Learning to Re-weight Domains for Continual Pre-training

### v1

@@ 3.1 Modeling the Heuristic Space with Trajectory Sampling
Start State Estimation
The start state $\rho_{s}$ can easily be determined when data from the source field are available. We randomly sample 1B tokens from the training data and utilize the same Nvidia domain classifier to organize the data into the defined domains. The start state is then estimated as the normalization of the sample numbers in each domain. In scenarios where the data from the source field are unavailable (Grattafiori et al., 2024; Liu et al., 2024a), we explore randomly sampled data from the target model as estimates for the start state. To prove the viability of this method, we experiment on five Pythia models (Biderman et al., 2023), as they are trained on the same open-source Pile dataset (Gao et al., 2020), where the ground-truth start state can be calculated. Specifically, we sample tokens simply by pre-pending the start-of-sentence token to start generation with a default temperature of $1.0$. The generated data are then passed through the domain classifier to estimate the start state. We calculate the KL divergence between the estimated start state and the ground-truth distribution obtained from the Pile dataset, and the results are presented in Fig. 3. As shown, sampling over 2,000 data points leads to a KL divergence of less than 0.1 on 4 out of 5 Pythia models. The estimated distribution converges on most models with over 3,000 samples. Larger models also show more accurate estimates and a faster convergence rate, possibly due to the higher quality of their generated data. The above results prove the effectiveness of using random samples from the target model as estimates for their start states.
@@ 20       $s_{d}\leftarrow 0$ ;
Data Mixing Trajectory Sampling
We randomly sample data mixing trajectories as the foundation of the training data for modeling the heuristic space. The random sampling process is based on the following principle:
The data mixing trajectories should be well-distributed across the defined action space, ensuring coverage of actions that both enhance and degrade model performance.
To ensure this principle, we design inductive scoring algorithms to rate each sampled distribution, guiding the next-action selection process. The detailed algorithm for the sampling process is provided in Algorithm 1. The function $CalculateInductiveScores$ describes the scoring algorithm. This function is designed based on three inductive biases that denote a potentially good distribution:
The target distribution is defined as the complement of the start state: probabilities for source-field domains are set to zero, while those for target-field domains are estimated based on their empirical distribution in the target field data. The target distribution encourages trajectories to gradually reduce reliance on source-field data and increase the coverage of target-field data, thereby accelerating the continual pre-training process.
Evaluation Environment Design and Feedback Collection
The evaluation environment is manually curated to assess model checkpoints. It is designed to be lightweight, yet accurately reflect target capabilities, providing effective supervision signals while minimizing computational overhead. Specifically, we select a small high-quality evaluation set $\mathcal{D}_{i}$ that well represents the $i$-th target field: $\{q_{j},r_{j}\}_{j=1}^{|\mathcal{D}_{i}|}$. For the model checkpoint $\mathcal{M}$, we compute the average per-token log probability on all question-answer pairs to reflect model performance on the $i$-th target field:
$\text{Score}(\mathcal{M},\mathcal{D}_{i})=\frac{1}{|\mathcal{D}_{i}|}\sum_{(q_{j},r_{j})\in\mathcal{D}_{i}}\frac{1}{|r_{j}|}\log P_{\mathcal{M}}(r_{j}\mid q_{j})$ (1)
$\text{reward}(\mathcal{M})=\left[\text{Score}(\mathcal{M},\mathcal{D}_{1}),\text{Score}(\mathcal{M},\mathcal{D}_{2}),...,\text{Score}(\mathcal{M},\mathcal{D}_{|\mathcal{D}|})\right]$ (2)
During implementation, the environment assesses the general capability of the checkpoints via the validation set of the MMLU dataset  [^2] https://huggingface.co/datasets/cais/mmlu with 1,531 high-quality general-domain questions and answers. The math reasoning capability is evaluated with 1,500 random samples from the training split of the MATH dataset  [^3] https://huggingface.co/datasets/EleutherAI/hendrycks_math. Leveraging this evaluation environment, we collect feedback data by training a small proxy model $\mathcal{M}_{p}$ with the LLaMA3 structure and 50M parameters on each sampled trajectory from scratch. The model checkpoint is evaluated on this environment at each data reweighting step, resulting in 27,266 feedbacks. Formally, for the $i$-th data mixing distribution $\rho_{i}\in\tau$, we obtain a tuple $(\rho_{i},reward(\mathcal{M}_{p}^{i}))$, where $reward(\mathcal{M}_{p}^{i})$ denotes the environment feedback for the model checkpoint $\mathcal{M}_{p}^{i}$ after training on the $i$-th domain reweighting step. Notably, the feedback at the start state is obtained with the initialized base proxy model.
3.2 Parameterize the Heuristic Space with Reinforcement Learning
We expect the sampled data mixing trajectories and the feedback from the environment to well represent the heuristic space for domain reweighting. We further parameterize these heuristics by training a model-based agent on these trajectories in a reinforcement learning-based paradigm.
Agent Model Structure
We determine the model structure for the data mixing agent with the following principles:
  • The structure should be designed to effectively model temporal sequences and support long-range interactions between data mixing distributions;
  • The data mixing agent should be lightweight to enable fast, low-cost inference and prevent unacceptable latency during target model training.
Based on the above principles, we utilize the Transformer (Vaswani et al., 2017) decoder architecture, which is widely applied to time series forecasting tasks (Zhang et al., 2024; Li et al., 2025) and facilitates long-range interactions between data points with its dot-product attention mechanism. To ensure fast inference, we stack two layers of Transformer, followed by a linear layer and Softmax to project the representations to the action space, with merely 2.1M parameters. Formally, at data reweighting step $t$, the agent $f$ predicts the domain distribution with the previous trajectory and environment feedback as follows:
$\rho_{t}=f(\tilde{\rho}_{0},\tilde{\rho}_{1},...,\tilde{\rho}_{t-1})$ (3)
$\tilde{{\rho}}_{i}=\left[\rho_{i};reward(\mathcal{M}_{i})\right],i=0,...,t-1$
where ; denotes concatenation, $\tilde{\rho}_{i}\in\mathcal{R}^{N+|\mathcal{D}|}$ denotes the input feature in the data reweighting step $i$, and $\rho_{t}$ denotes the agent’s output action at step $t$.
SFT-based Warming Up
We first perform Supervised Fine-Tuning (SFT) to reduce the parameter searching space in the reinforcement learning phase. We train the agent from scratch on the high-quality $\mathcal{T}_{top1}$ trajectories with a simple MSE loss. At the data reweighting step $t$, the agent is optimized as follows:
Off-policy Optimization with Conservative Q-Learning
Based on the warmed-up agent model, we further parameterize the heuristic space via reinforcement learning, where the algorithm selection is based on the following two principles:
  • The agent model is trained in an offline, off-policy setting using data collected from proxy models, without access to the evaluation environment during training;
  • The agent’s actions are probability distributions sampled from a continuous domain space.
Following the above principles, we select Conservative Q-Learning (CQL) (Kumar et al., 2020) as the optimization algorithm. CQL prevents overestimation of Q-values for out-of-distribution actions by encouraging the learned Q-function to be conservative. Specifically, CQL introduces a conservative penalty for the Q-function optimization process:
$\mathcal{L}_{\text{CQL}}(Q)=$ $\underbrace{\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q(s,a)-\left(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})\right)\right)^{2}\right]}_{\text{Bellman error}}$ (5)
$+\alpha\cdot\underbrace{\left(\mathbb{E}_{s\sim\mathcal{D}}\left[\log\sum_{a}\exp(Q(s,a))\right]-\mathbb{E}_{(s,a)\sim\mathcal{D}}[Q(s,a)]\right)}_{\text{Conservative penalty}}$
CQL is then trained in an actor-critic (Sutton et al., 1999) structure, where the data agent acts as the actor model, and another neural network is initialized from scratch as the critic model (Q-function).
During implementation, we randomly sample fragments $\tau$ (don’t have to be full trajectories) from the data mixing trajectory set $\mathcal{T}$. At domain re-weighting step $t$, $s=[\rho_{0},\rho_{1},...,\rho_{t-1}]$, $a=\rho_{t}$, and $s^{\prime}=[\rho_{0},\rho_{1},...,\rho_{t}]$. The scalar reward value $r$ is obtained as the gain of a linear combination of environment feedback $reward(\mathcal{M}_{p}^{t})$ compared to that of the last step:
$r=\sum_{i=1}^{|D|}\lambda_{i}Score(\mathcal{M}_{p}^{t},D_{i})-\sum_{i=1}^{|D|}\lambda_{i}Score(\mathcal{M}_{p}^{t-1},D_{i})$ (6)
During implementation, we set all coefficients to be equal: $\lambda_{i}=\frac{1}{|D|}$. The critic model $f^{\prime}$ is parameterized by another single-layer Transformer decoder, followed by a linear layer and sigmoid function to project the representations into a Q-value scalar, with the following inference function:
$Q(s,a)=f^{\prime}(\rho_{0},\rho_{1},...,\rho_{t})$ (7)
The actor and critic models are trained in this function until convergence.
@@ 3.2 Parameterize the Heuristic Space with Reinforcement Learning
Input: $N$ domains of source data $\{S_{1},S_{2},...,S_{N}\}$ and target data $\{T_{1},T_{2},...,T_{N}\}$, the agent $f$, Max data reweighting steps $M_{tgt}$, Reweighting sample number per step $R_{tgt}$, the target model $\mathcal{M}_{tgt}$, the evluation environment $\mathcal{E}$.
Output: The continually pretrained target model checkpoint ^$\hat{\mathcal{M}}_{tgt}$
@@
^$\hat{\mathcal{M}}_{tgt}\leftarrow\mathcal{M}_{tgt}$ ;
@@
       $reward(\hat{\mathcal{M}}_{tgt})\leftarrow$ GetEnvFeedback(^$\hat{\mathcal{M}}_{tgt}$, $\mathcal{E}$) ;
        // Get environment feedback for the current model checkpoint.
@@
        // Concatenate each trajectory with the corresponding feedback.
@@
        // Obtain the data reweighting distribution from the data mixing agent.
@@
      $\mathcal{B}_{t}\leftarrow$ sample($\{S_{i}\}_{1}^{N}$,$\{T_{i}\}_{1}^{N}$,$\rho_{t}$);
        // Sample a batch from the domain data according to the current distribution.
@@
11      Update weights for ^$\hat{\mathcal{M}}_{tgt}$ with the training loss (^$\mathcal{L}(\hat{\mathcal{M}}_{tgt},\mathcal{B}_{t})$;
@@ 13      Append $\rho_{t}$ to $\mathcal{T}_{tgt}$;
3.3 Domain Reweighting with Data Mixing Agent
The mechanism of the domain reweighting process with Data Mixing Agent is described in Algorithm 2. During continual pre-training, the agent directly determines the distribution for the next domain re-weighting step on the fly, considering the previous states in the data mixing trajectory and the corresponding environment feedback. This MDP continues until the target data is fully leveraged or a predetermined computation budget is reached. We expect the agent to optimally curate the training recipe by balancing performance across all target fields while minimizing the use of source-field data tokens to reduce computational cost. We also expect the agent’s learned heuristics to generalize to unseen target models, data mixtures, and even target domains. This generalization is crucial to avoid repeated trajectory sampling and agent retraining when adapting to new continual pre-training scenarios, thereby significantly reducing overall computational cost.
4 Experiments
In this section, we evaluate the performance of Data Mixing Agent in improving the math reasoning and code generation capabilities of target models via continual pretraining. We comprehensively compare the performance of the agent to strong baseline methods across 4 target models on 8 general benchmarks, 4 math reasoning benchmarks, and 2 code generation benchmarks.
4.1 Experimental Settings
Target Models
@@ 4.1 Experimental Settings
We aim to rigorously evaluate domain reweighting methods on target models that do not possess math or coding capabilities. Since most open-source models have been optimized on large-scale data from the math reasoning or code generation field, we pre-train three models from scratch, with the same LLaMA3 model architecture (Grattafiori et al., 2024) of 32 Transformer layers and 3B model parameters, on 100B randomly sampled tokens from the DCLM (Li et al., 2024), Fineweb-Edu (Penedo et al., 2024), and Nemotron-CC (Su et al., 2024) dataset, resulting in three target models: LLaMA-3B-DCLM-100B, LLaMA-3B-FWE-100B, LLaMA-3B-Nemotron-100B. We also include the Pythia-1.4B model (Biderman et al., 2023) to evaluate performance on existing open-source models and scenarios when data from the source field is not directly available.
Baseline Methods
@@
  • Base Model: direct evaluation of the target models on the benchmarks, reflecting model capabilities before the continual pertaining phase;
@@
  • RegMix (Liu et al., 2024b): one of the state-of-the-art domain re-weighting methods. It trains large quantities of 1B-sized small proxy models (512 models in our implementation) on random domain distributions, then evaluates these models on the target benchmarks. The best data mixing recipe is determined by fitting a regression model to the feedback and selecting distributions that lead to the highest scores;
@@
Target Model Training Data

### v2

@@ Appendix A Agent Model
Appendix B Start State Estimation with Synthetic Data
In scenarios where the data from the source field are unavailable (Grattafiori et al., 2024; Liu et al., 2024a), we explore randomly sampled data from the target model as estimates for the start state. To prove the viability of this method, we experiment on five Pythia models (Biderman et al., 2023), as they are trained on the same open source Pile dataset (Gao et al., 2020), where the ground-truth start state can be calculated using the same method as in the paper. Specifically, we first randomly sample tokens from the target model simply by pre-pending the start-of-sentence token to start generation with a default temperature of $1.0$. The generated data are then passed through the same Nvidia domain classifier to obtain a domain label. The estimated start state is calculated by normalizing sample numbers on each domain in the generated data.
To prove the viability of such estimation, we calculate the KL divergence between the estimated start state and the ground-truth distribution obtained from the Pile dataset, and the results are presented in Fig. 4. As shown, sampling over 2,000 random samples already leads to a KL divergence of less than 0.1 on 4 out of 5 Pythia models, showing a fast convergence rate of the proposed estimation method. The estimated distribution also converges on most models with over 4,000 samples. In addition, larger models such as Pythia-6.9B also show more accurate estimates and a faster convergence rate, possibly due to the higher quality of their generated data, leading to more accurate domain classification. The above results prove the effectiveness of using random samples from the target model as estimates for their start states. This method is further utilized in experiments on the Pythia-1.4B model.
Appendix C Implementation Details
In this section, we provide more technical implementation details of Data Mixing Agent, including the trajectory sampling algorithm, the evaluation environment feedback design, the data mixing agent training process, and our evaluation process on various settings.
@@ 24 $s_{d}\leftarrow 0$ ;
Input: NN domains of source data $\{S_{1},S_{2},...,S_{N}\}$ and target data $\{T_{1},T_{2},...,T_{N}\}$, the agent $f$, Max data reweighting steps $M_{tgt}$, Reweighting sample number per step $R_{tgt}$, the target model $\mathcal{M}_{tgt}$, the evluation environment $\mathcal{E}$.
Output: The continually pretrained target model checkpoint $\hat{\mathcal{M}}_{tgt}$
@@
$\hat{\mathcal{M}}_{tgt}\leftarrow\mathcal{M}_{tgt}$ ;
@@
 $reward(\hat{\mathcal{M}}_{tgt})\leftarrow$ GetEnvFeedback($\hat{\mathcal{M}}_{tgt}$, $\mathcal{E}$) ;
 // Get environment feedback.
@@
 // Concat trajectories with the corresponding feedback.
@@
 // Obtain domain reweighting distribution from the agent.
@@
 $\mathcal{B}_{t}\leftarrow$ sample($\{S_{i}\}_{1}^{N}$$\{T_{i}\}_{1}^{N}$$\rho_{t}$);
 // Sample domain data with the current distribution.
@@
11  Update weights for $\hat{\mathcal{M}}_{tgt}$ with the training loss $\mathcal{L}(\hat{\mathcal{M}}_{tgt},\mathcal{B}_{t})$;
@@ 13  Append $\rho_{t}$ to $\mathcal{T}_{tgt}$;
C.1 Trajectory Sampling
Algorithm
The detailed algorithm for the trajectory sampling process is provided in Algorithm 1. The function $CalculateInductiveScores$ describes the scoring algorithm. This function is designed based on three inductive biases that denote a potentially good distribution:
The target distribution is defined as the complement of the start state: probabilities for source-field domains are set to zero, while those for target-field domains are estimated based on their empirical distribution in the target field data. This target distribution serves as a soft constraint that encourages trajectories to gradually reduce reliance on source-field data and increase the coverage of target-field data, thereby accelerating the continual pre-training process and saving computation costs.
Implementation
C.2 Evaluation Environment Design and Feedback Collection
During implementation, the environment assesses the general capability of the checkpoints via 500 high-quality general-domain questions and answers from the validation set of the MMLU dataset  [^2] https://huggingface.co/datasets/cais/mmlu. The math reasoning capability is evaluated with 500 random samples from the training split of the MATH dataset  [^3] https://huggingface.co/datasets/EleutherAI/hendrycks_math.
Leveraging this evaluation environment, we collect feedback data by training a small proxy model $\mathcal{M}_{p}$ with the LLaMA model structure and 50M parameters on each sampled trajectory from scratch. The model checkpoint is evaluated on this environment at each data reweighting step, resulting in 27,266 feedbacks. Formally, for the $i$-th data mixing distribution $\rho_{i}\in\tau$, we obtain a tuple $(\rho_{i},\mathcal{R}(\mathcal{M}_{p}^{i}))$, where $\mathcal{R}(\mathcal{M}_{p}^{i})$ denotes the environment feedback for the model checkpoint $\mathcal{M}_{p}^{i}$ after training on the $i$-th domain reweighting step. Notably, the feedback at the start state is obtained with the initialized base proxy model.
C.3 SFT-based Warming Up
We first perform Supervised Fine-Tuning (SFT) to reduce the parameter searching space in the reinforcement learning phase. We train the agent from scratch on the high-quality $\mathcal{T}_{top1}$ trajectory subset (obtained during the trajectory sampling process) with a simple MSE loss. At the data reweighting step $t$, the agent is optimized as follows:
C.4 Conservative Q-Learning
We select Conservative Q-Learning (CQL) (Kumar et al., 2020) as the optimization algorithm. CQL prevents overestimation of Q-values for out-of-distribution actions by encouraging the learned Q-function to be conservative. Specifically, CQL introduces a conservative penalty for the Q-function optimization process with the following training objective:
$\underbrace{\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q(s,a)-\left(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})\right)\right)^{2}\right]}_{\text{Bellman error}}$ (5)
$+\alpha\cdot\underbrace{\left(\mathbb{E}_{s\sim\mathcal{D}}\left[\mathbb{E}_{a^{\prime}\sim\mathcal{U(\mathcal{A})}}(Q(s,a^{\prime}))\right]-\mathbb{E}_{(s,a)\sim\mathcal{D}}[Q(s,a)]\right)}_{\text{Conservative penalty}}$
where the first term denotes the standard Bellman error for Q-learning, and the second term denotes a conservative penalty to minimize the Q value expectations for randomly sampled actions from the current state under another distribution $\mathcal{U}(\mathcal{A})$. The target model and the Q function are then trained in an actor-critic (Sutton et al., 1999) structure, where the data agent acts as the actor model, optimized with policy gradient to enhance actions that maximize the Q value feedback from the critic model. Another neural network is initialized from scratch as the parameterized Q function, which acts as the critic model to evaluate model actions and optimized via Eqn. 5.
During implementation, we randomly sample fragments $\tau$ (don’t have to be full trajectories) from the data mixing trajectory set $\mathcal{T}$. At domain re-weighting step $t$, $s=[\rho_{0},\rho_{1},...,\rho_{t-1}]$, $a=\rho_{t}$, and $s^{\prime}=[\rho_{0},\rho_{1},...,\rho_{t}]$. The scalar reward value $r$ is obtained as the gain of a linear combination of environment feedback $\mathcal{R}(\mathcal{M}_{p}^{t})$ compared to that of the last step:
$r=\sum_{i=1}^{|D|}\lambda_{i}\mathcal{S}(\mathcal{M}_{p}^{t},D_{i})-\sum_{i=1}^{|D|}\lambda_{i}\mathcal{S}(\mathcal{M}_{p}^{t-1},D_{i})$ (6)
During implementation, we set all coefficients to be equal to encourage balanced consideration of target field capabilities: $\lambda_{i}=\frac{1}{|D|}$. The critic model $f^{\prime}$ is parameterized by another single-layer Transformer decoder, followed by a linear layer and sigmoid function to project the representations into a Q-value scalar, with the following inference function:
$Q(s,a)=f^{\prime}(\rho_{0},\rho_{1},...,\rho_{t})$ (7)
The agent model and the Q-function model are iteratively optimized in this actor-critic manner until convergence.
The advantage of CQL over SFT is expected. The theoretical foundation of the data mixing agent lies in learning from both actions that enhance and degrade model performance. CQL, as SOTA off-policy RL algorithm, is inherently designed to optimize policies using both positive and negative signals from sampled trajectories. Specifically, it promotes actions associated with high Q-values while suppressing those associated with low Q-values.
In contrast, SFT is an instance of imitation learning, which optimizes the policy by fitting it to high-quality (i.e., “good”) trajectories. By design, SFT cannot exploit informative signals from suboptimal or “bad” trajectories, where the policy should instead learn to avoid certain actions. When incorporated, such trajectories may even introduce noise rather than useful learning signals. For this reason, we employ SFT only as a warm-up strategy and subsequently apply CQL to fully optimize the policy over all trajectories.
We believe the strong generalization capability of the agent stems from the same intuition that motivated the development of the data mixing agent in the first place: there exists a rich heuristic space for domain reweighting. These heuristics are largely model- and data-agnostic, and can therefore be unified within a compact agent model to guide data mixing trajectories in a principled manner. Our approach differs from prior work in that, rather than relying on manually designed algorithms, we train a model to parameterize these heuristics and enable automated domain reweighting. As a result, we expect our method to achieve comparable or superior generalization performance, consistent with the empirical success of previous heuristic-based approaches.
C.5 System Pipeline of Data Mixing Agent
Following Google’s definition  [^4] https://cloud.google.com/discover/what-are-ai-agents, an AI agent is a software system that autonomously pursues goals and completes tasks by exhibiting reasoning, planning, and adaptation capabilities. Under this definition, the proposed data mixing agent qualifies as an AI agent. Concretely, it is implemented as a Transformer-based learning system that operates as an independent decision-making module. Its objective is to predict an optimal data mixing recipe for the next domain reweighting step, given the historical training trajectory as input. By encoding past data mixing decisions and their associated outcomes, the agent effectively maintains memory over previous trajectories and reasons over them to plan future data allocation strategies. Moreover, empirical results demonstrate that the agent can learn and adapt across domains, generalizing effectively to diverse target fields such as math reasoning and code generation.
In the overall system pipeline, the data mixing agent is invoked directly at each reweighting step to predict the next set of domain mixing weights, rather than calling a language model for decision making. This design aligns with the general notion of AI agents as decision-making systems that are not necessarily language models themselves.
Appendix D Experimental Settings
D.1 Target Models
@@
We aim to rigorously evaluate domain reweighting methods on target models that do not possess math or coding capabilities. Since most open-source models have been optimized on large-scale data from the math reasoning or code generation field, we pre-train three models from scratch, with the same LLaMA3 model architecture (Grattafiori et al., 2024) of 32 Transformer layers and 3B model parameters, on 100B randomly sampled tokens from the DCLM (Li et al., 2024), Fineweb-Edu (Penedo et al., 2024), and Nemotron-CC (Su et al., 2024) dataset, resulting in three target models: LLaMA-DCLM, LLaMA-FWE, LLaMA-Nemo. We also include the Pythia-1.4B model (Biderman et al., 2023) to evaluate performance on existing open-source models and scenarios when data from the source field is not directly available.
D.2 Baseline Methods
@@
  • Base Model: direct evaluation of the target models on the benchmarks, reflecting model capabilities before the continual pretraining phase;
@@
  • RegMix (Liu et al., 2024b): state-of-the-art static domain re-weighting method. It trains large quantities of 1B-sized small proxy models with LLaMA structure (512 models in our implementation) on random domain distributions, then evaluates these models on the target benchmarks. The best data mixing recipe is determined by fitting a regression model to the feedback and selecting distributions that lead to the highest scores. We mostly follow the settings of Wettig et al. (2025) in implementing the RegMix algorithm;
  • Dynamic Batch Loading (DBL) (Xia et al., 2023): State-of-the-art dynamic domain reweighting method that modifies the data mixture on the fly. It first estimates a reference optimal loss on each domain reweighting step via the scaling law (Hoffmann et al., 2022), then computes the excess loss between the current evaluation loss and the optimal loss on each domain, increasing data ratios for slowly learning domains for the next training data batch. We skip the procedure for fitting the scaling function and directly utilize the loss of the base model on the general environment set as the reference loss for the general domain, the loss of the naively trained model on the math environment set as the reference loss for math reasoning, and the loss of the naively trained model on the code environment set as the reference loss for code generation. The DBL uses the same reweighting samples per step as Data Mixing Agent until the target field data is fully covered. We only use DBL as a baseline on the 2-dimensional domain space, due to the lack of evaluation sets for all domains in the 52-dimensional domain space;
@@
D.3 Target Model and Data